# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v2

@@ Abstract
In the medical field, the limited availability of large-scale datasets and labor-intensive annotation processes hinder the performance of deep models. Diffusion-based generative augmentation approaches present a promising solution to this issue, having been proven effective in advancing downstream medical recognition tasks. Nevertheless, existing works lack sufficient semantic and sequential steerability for challenging video/3D sequence generation, and neglect quality control of noisy synthesized samples, resulting in unreliable synthetic databases and severely limiting the performance of downstream tasks. In this work, we present Ctrl-GenAug, a novel and general generative augmentation framework that enables highly semantic- and sequential-customized sequence synthesis and suppresses incorrectly synthesized samples, to aid medical sequence classification. Specifically, we first design a multimodal conditions-guided sequence generator for controllably synthesizing diagnosis-promotive samples. A sequential augmentation module is integrated to enhance the temporal/stereoscopic coherence of generated samples. Then, we propose a noisy synthetic data filter to suppress unreliable cases at the semantic and sequential levels. Extensive experiments on 3 medical datasets, using 11 networks trained on 3 paradigms, comprehensively analyze the effectiveness and generality of Ctrl-GenAug, particularly in underrepresented high-risk populations and out-domain conditions.
keywords:
Classification, Controllable Generative Augmentation, Medical Sequence Synthesis, Diffusion Model
journal: ***
@@ 1 Introduction
Dynamic information in medical imaging (e.g., video sequences) plays a vital role in clinical diagnosis. Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases (Zhou et al., 2023; Wang et al., 2024c). Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers (Varoquaux and Cheplygina, 2022; Huang et al., 2022). This risk presents a significant obstacle to deploying models in real-world medical settings. To address these issues, generative augmentation paradigms have employed advanced generative models to synthesize medical samples, thereby augmenting relevant diagnostic tasks.
@@
Figure 1: Datasets description: (a) Carotid US videos with various stenosis gradings. Red and yellow boxes represent plaques and residual lumens, respectively. (b) Thyroid US videos with different TI-RADS levels, where blue boxes indicate nodules. (c) Cardiac MRI volumes with distinct diseases. Three key anatomical structures associated with diagnosis are highlighted with masks, including the left ventricle (blue), myocardium (green), and right ventricle (red).
Intuitively, building customized, high-fidelity, and sequentially coherent synthetic medical databases, and effectively utilizing them, is essential for enhancing sequence recognition. However, this task presents several challenges. First, lesions (Fig. 1(a-b)) or structures (Fig. 1(c)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc). This may seriously confuse diffusion learning, thus causing uncontrollable and unreliable generation. Second, artifacts and noises (e.g., US speckle noise) may prevent models from accurately perceiving vital anatomical targets, hindering the synthesis of high-fidelity sequences. Third, the complex dynamic changes (Fig. 1(a-b)) and the varying sizes of anatomical regions (Fig. 1(c)) challenge the coherence of synthesized sequences. The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning (He et al., 2023). Lastly, even a well-designed sequence generator cannot always guarantee high-quality synthesis due to random sampling, and unsatisfactory synthetic samples may negatively impact subsequent classifier learning. In summary, from model- and data-centric perspectives, the challenges of the task can be concluded as: 1) How can we design the generator architecture to achieve satisfactory sequence synthesis? 2) How can we discriminatively filter out potentially harmful synthetic samples?
In this work, we propose a novel controllable generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification tasks on different organs and modalities. Concretely, Ctrl-GenAug synthesizes customized, high-quality sequences using conditional diffusion models and applies quality control to the generated sequences, to supplement medical datasets and improve classification. This is a general method for enhancing medical sequence classification. We believe this is the first comprehensive study to analyze the impact of the controllable generative augmentation scheme on promoting multi-modal medical sequence classification. Our contributions are three-fold:
  • 1.
@@
  • 3.
    Comprehensive experiments on 3 medical datasets, with 11 downstream networks trained on 3 paradigms, show that our proposed framework can improve diagnostic performance, especially within underrepresented high-risk sets. Ctrl-GenAug also increases out-domain robustness, further indicating its practicality in clinical scenarios.
@@ 2.1 Controllable Video Synthesis with Diffusion Models
Diffusion models, a class of generative models, have recently attracted significant attention, especially text-to-video (T2V) synthesis (Blattmann et al., 2023; Ho et al., 2022a; Singer et al., 2022; Zhou et al., 2022). However, these methods did not achieve precise control over the visual appearance and dynamics of the synthesized video, resulting in limited practicality. To mitigate this issue, researchers have concentrated on incorporating multimodal conditions into T2V frameworks to guide controllable synthesis. Specifically, AnimateDiff (Guo et al., 2023) and MoonShot (Zhang et al., 2024a) introduced text and image conditions to control the visual appearance. Most recently, several studies (Zhang et al., 2023; Zhao et al., 2023; Wang et al., 2024b) achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences). Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers (Zhang et al., 2023) and S2 condition encoder (Wang et al., 2024b)) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction. Besides, dense optical flow (Zach et al., 2007) was adopted to control sequential consistency (Liang et al., 2023; Ni et al., 2023). The above methods have shown promising controllability in natural video synthesis. However, for medical synthesis, acquiring abundant dense annotations (e.g., mask sequences) as control signals is impractical and unaffordable. Additionally, medical videos have unique attributes compared with natural ones, such as blurred anatomical regions and complex structure variations. Hence, they may not suit the medical video synthesis tasks.
In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame (Reynaud et al., 2023), ejection fractions (Reynaud et al., 2023), single semantic mask (Van Phi et al., 2024; Nguyen et al., 2024), and single sketch image (Zhou et al., 2024a). Zhou et al. (Zhou et al., 2024a) additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms. Besides, Li et al. (Li et al., 2024a) proposed to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors. However, the above methods may have the following drawbacks: unsteerable semantics due to the lack of attribute guidance (e.g., category or target shape), poor control over sequences, etc. Thus, they may be unsuitable for direct adoption in the generative augmentation tasks.
Figure 2: t-SNE visualization of features of synthetic and real training data by a pre-trained I3D (Carreira and Zisserman, 2017) model on our three datasets.
@@ 2.2 Generative Augmentation with Diffusion Models
Synthetic data can provide intrinsically diverse characteristics and visual appearances (Luo et al., 2024; Pan et al., 2023), thus enhancing model learning. This aligns with the observation shown in Fig. 2. Previous works have succeeded in utilizing synthetic data by diffusion models to enhance natural image classification (He et al., 2023; Sarıyıldız et al., 2023; Trabucco et al., 2023; Azizi et al., 2023), segmentation (Sankaranarayanan et al., 2018; Wu et al., 2023b; Yang et al., 2024), and detection (Feng et al., 2024). Recently, Singh et al. (Singh et al., 2024) performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks (Shang et al., 2023; Farooq et al., 2024; Ktena et al., 2024; Zhang et al., 2024c; Dorjsembe et al., 2024b; Li et al., 2024b). Luo et al. (Luo et al., 2024) marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks. Nevertheless, current studies have two main limitations:
@@ 3.1 Model Overview
Fig. 3 shows the pipeline of using our proposed controllable generative augmentation framework to aid medical sequence classification. First, we propose a sequence generator that enables perceiving multiple semantic and sequential conditions to guide controllable and diverse generation. Then, an efficient noisy data filter is introduced to suppress unsatisfactory synthetic sequences. Last, the quality-controlled synthetic sequences and real ones will work together to improve the performance of arbitrary classifiers. To formulate customized, high-fidelity, and coherent synthetic databases to boost classification, we design the whole framework from model-centric, i.e., sequence generator (Secs. 3.23.3, and 3.4) and data-centric, i.e., data filter (Sec. 3.5) perspectives.
@@ 3.2 Basic Architecture of Sequence Generator
Fig. 4 shows the pipeline of our proposed sequence generator. It supports customized and high-quality medical sequence generation via multimodal conditions, using a two-stage training scheme. In the pretraining stage, it attends to high-fidelity visual features learning for controllable image synthesis. While in the finetuning stage, the domain-specific visual knowledge acquired from the previous stage is reused and focuses on sequential patterns modeling for customized sequence synthesis. During inference, given a single or combination of multimodal conditions as control signal inputs, high-quality and steerable sequence generation from Gaussian noise can work. We then provide a short background of video diffusion models in Sec. 3.2.1, followed by a detailed description of the basic architecture of our generator in Secs. 3.2.2 and 3.2.3.
3.2.1 Preliminaries of Video DDPMs
Video DDPMs aim to learn complex distributions by iteratively denoising corrupted inputs using diffusion methods (Sohl-Dickstein et al., 2015; Ho et al., 2020). Given a video $\mathit{x}_{0}\sim\mathit{q}$ and timestep $\mathit{T}$, DDPMs first produce a sequence of noisy inputs via diffusion process $\mathit{q}(\mathit{x}_{t}|\mathit{x}_{0},\mathit{t}),\mathit{t}\in{1,2,...,\mathit{T}}$, which progressively adds Gaussian noise $\epsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$ to $\mathit{x}_{0}$. To ease the computational burden of traditional pixel-space training, video latent diffusion models (Blattmann et al., 2023) perform the diffusion process in latent space of a variational autoencoder (Esser et al., 2021) (VAE). The model is then trained to estimate the parameterized Gaussian transition $\mathit{p}(\mathit{x}_{t-1}|\mathit{x}_{t})$ via a denoising network $\theta$. Mathematically, the optimized objective can be a simplified variant of the variational lower bound:
$\begin{split}\min_{\theta}\mathbb{E}_{\mathit{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},\mathit{t}}||\epsilon-\epsilon_{\theta}(\mathit{z}_{t},\mathbf{c},\mathit{t})||^{2},\end{split}$ (1)
@@ 3.2.1 Preliminaries of Video DDPMs
3.2.2 Factorized Learning of Visual Features and Sequential Patterns
Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis (Blattmann et al., 2023). However, due to the scarcity and complexity of medical data, learning visual features and sequential patterns at the same time is very challenging and imposes high training costs. To this end, we propose to factorize both learning by first pretraining a latent diffusion model (LDM) and then finetuning its sequence counterpart. In this way, the generator enables realistic sequence synthesis, while easing the training burden.
3.2.3 2D-to-3D Model Inflation
During pretraining, we develop a 2D UNet (Ronneberger et al., 2015) in LDM to predict noises for image synthesis. Referring to (Ho et al., 2022b), we extend the 2D UNet to a 3D version via an inflation scheme to build the sequence LDM. Specifically, all the spatial convolution layers are inflated to pseudo-3D counterparts by expanding the kernels at the sequential dimension (e.g., $3\times 3\rightarrow 1\times 3\times 3$ kernel). Besides, we perform sequential insertion by adding sequential attention (SA) layers (see Fig. 4). Fig. 5(a) shows the principle of the SA mechanism, where each patch queries to those at the same spatial position and across frames/slices. This design encourages the generator to model the sequential patterns, while not significantly altering visual feature distribution baked in the LDM (Zhang et al., 2024a). Overall, sequence LDM enables inheriting the rich visual concepts preserved in LDM and focusing on sequential pattern aggregation, making the model learning efficient.
Figure 4: Pipeline of our proposed sequence generator. The MedSAM image encoder (Huang et al., 2024b) is used for domain-specific image prior feature extraction.

### v3

@@ Abstract
In the medical field, the limited availability of large-scale datasets and labor-intensive annotation processes hinder the performance of deep models. Diffusion-based generative augmentation approaches present a promising solution to this issue, having been proven effective in advancing downstream medical recognition tasks. Nevertheless, existing works lack sufficient semantic and sequential steerability for challenging video/3D sequence generation, and neglect quality control of noisy synthesized samples, resulting in unreliable synthetic databases and severely limiting the performance of downstream tasks. In this work, we present Ctrl-GenAug, a novel and general generative augmentation framework that enables highly semantic- and sequential-customized sequence synthesis and suppresses incorrectly synthesized samples, to aid medical sequence classification. Specifically, we first design a multimodal conditions-guided sequence generator for controllably synthesizing diagnosis-promotive samples. A sequential augmentation module is integrated to enhance the temporal/stereoscopic coherence of generated samples. Then, we propose a noisy synthetic data filter to suppress unreliable cases at the semantic and sequential levels. Extensive experiments on 5 medical datasets with 4 different modalities, including comparisons with 15 augmentation methods and evaluations using 11 networks trained on 3 paradigms, comprehensively demonstrate the effectiveness and generality of Ctrl-GenAug, particularly with pronounced performance gains in underrepresented high-risk populations and out-domain conditions. Codes, models, and synthetic databases are available at https://github.com/XinRuiZhou0106/Ctrl-GenAug.
@@ 1 Introduction
Dynamic information in medical imaging (e.g., video sequences) plays a vital role in clinical diagnosis. Recent deep learning-based classifiers have shown the ability to improve the diagnostic accuracy of different diseases (Zhou et al., 2023; Wang et al., 2024b). Despite rapid advancements, the performance of current advanced solutions is still limited by several issues: 1) The scarcity of dynamic clinical cases, coupled with the high annotation cost, limits data availability; 2) Imbalanced data distribution, driven by the rarity of high-risk positive cases, skews model training; 3) Deep models are prone to brittle performance degradation when tested on out-domain data, e.g., cases from different medical centers (Varoquaux and Cheplygina, 2022; Huang et al., 2022). This risk presents a significant obstacle to deploying models in real-world medical settings. To address these issues, generative augmentation paradigms have employed advanced generative models to synthesize medical samples, thereby augmenting relevant diagnostic tasks.
@@
Figure 1: Datasets description: (a) Carotid US videos with various stenosis gradings. Red and yellow boxes represent plaques and residual lumens, respectively. (b) Thyroid US videos with different TI-RADS levels, where blue boxes indicate nodules. (c) Cardiac Cine-MRI volumes with distinct diseases. Three key anatomical structures associated with diagnosis are highlighted with masks, including the left ventricle (blue), myocardium (green), and right ventricle (red). (d) Lung CT volumes exhibiting varying degrees of tissue abnormalities with COVID-19. (e) Knee T2-weighted MRI volumes with and without anterior cruciate ligament (ACL) tears.
Intuitively, building customized, high-fidelity, and sequentially coherent synthetic medical databases, and effectively utilizing them, is essential for enhancing sequence recognition. However, this task presents several challenges. First, lesions (Fig. 1(a-b)) or structures (Fig. 1(c-e)) of the same disease category exhibit large visual variances (e.g., shapes, intensities, positions, etc). This may seriously confuse diffusion learning, thus causing uncontrollable and unreliable generation. Second, artifacts and noises (e.g., US speckle noise) may prevent models from accurately perceiving vital anatomical targets, hindering the synthesis of high-fidelity sequences. Third, the complex dynamic changes (Fig. 1(a-b)) and the varying sizes of anatomical regions (Fig. 1(c-e)) challenge the coherence of synthesized sequences. The fourth challenge lies in the domain gap between synthetic and real samples, where achieving real-domain customization is crucial for effectively utilizing synthetic samples in downstream learning (He et al., 2023). Lastly, even a well-designed sequence generator cannot always guarantee high-quality synthesis due to random sampling, and unsatisfactory synthetic samples may negatively impact subsequent classifier learning. In summary, from model- and data-centric perspectives, the challenges of the task can be concluded as: 1) How can we design the generator architecture to achieve satisfactory sequence synthesis? 2) How can we discriminatively filter out potentially harmful synthetic samples?
In this work, we propose a novel controllable generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification tasks on different organs and modalities. Concretely, Ctrl-GenAug synthesizes customized, high-quality sequences using conditional diffusion models and applies quality control to the generated sequences, to supplement medical datasets and improve classification. This is a general method for enhancing medical sequence classification. We believe this is the first comprehensive study to analyze the impact of controllable generative augmentation on multi-modal medical sequence classification. Our contributions are three-fold:
@@
  • Comprehensive experiments on 5 medical datasets with 4 modalities (US, CT, Cine/T2-weighted MRI), including benchmarking against 15 augmentation methods and evaluations using 11 downstream classifiers trained on 3 paradigms, show that our approach consistently achieves superior performance across diverse diagnostic tasks, while maintaining robust effectiveness with multiple classifier architectures. Ctrl-GenAug also improves diagnostic performance in underrepresented high-risk sets and enhances out-domain robustness, underscoring its practicality in real clinical scenarios.
@@ 2.1 Controllable Video Synthesis with Diffusion Models
Diffusion models, a class of generative models, have recently attracted significant attention, especially text-to-video (T2V) synthesis (Blattmann et al., 2023; Ho et al., 2022a; Singer et al., 2022; Zhou et al., 2022). However, these methods did not achieve precise control over the visual appearance and dynamics of the synthesized video, resulting in limited practicality. To mitigate this issue, researchers have concentrated on incorporating multimodal conditions into T2V frameworks to guide controllable synthesis. Specifically, AnimateDiff (Guo et al., 2023) and MoonShot (Zhang et al., 2024a) introduced text and image conditions to control the visual appearance. Most recently, several studies (Zhang et al., 2023; Zhao et al., 2023; Wang et al., 2023) achieved fine-grained spatial compositional control by integrating additional diverse conditions (i.e., sketch, mask, and depth sequences). Meanwhile, two spatial-sequential (S2) modeling strategies (fully attention layers (Zhang et al., 2023) and S2 condition encoder (Wang et al., 2023)) were introduced to capture the dynamics of sequential conditions, promoting the inter-frame features interaction. Besides, dense optical flow (Zach et al., 2007) was adopted to control sequential consistency (Liang et al., 2023; Ni et al., 2023). The above methods have shown promising controllability in natural video synthesis. However, for medical synthesis, acquiring abundant dense annotations (e.g., mask sequences) as control signals is impractical and unaffordable. Additionally, medical videos have unique attributes compared with natural ones, such as blurred anatomical regions and complex structure variations. Hence, they may not suit the medical video synthesis tasks.
Figure 2: t-SNE visualization of features of synthetic and real training data by a pre-trained I3D (Carreira and Zisserman, 2017) model on our five datasets. Our Ctrl-GenAug produces synthetic samples with diverse feature patterns that enrich the training distribution.
In the medical field, several diffusion-based methods have been proposed to generate photorealistic echocardiography videos using the conditional guidance of a single frame (Reynaud et al., 2023), ejection fractions (Reynaud et al., 2023), single semantic mask (Van Phi et al., 2024; Nguyen et al., 2024), and single sketch image (Zhou et al., 2024a). Zhou et al. (2024a) additionally leveraged mitral valve (MV) skeletons to control the complex motion trajectories of MV in the generated echocardiograms. Besides, Li et al. (2024a) proposed to generate endoscopy videos that simulate clinical scenes by integrating a spatial-temporal transformer with 2D vision foundation model priors. However, the above methods may have the following drawbacks: unsteerable semantics due to the lack of attribute guidance (e.g., category or target shape), poor control over sequences, etc. Thus, they may be unsuitable for direct adoption in the generative augmentation tasks.
@@ 2.2 Generative Augmentation with Diffusion Models
Synthetic data can provide intrinsically diverse characteristics and visual appearances (Luo et al., 2024; Pan et al., 2023), thus enhancing model learning. This aligns with the observation shown in Fig. 2. Previous works have succeeded in utilizing synthetic data by diffusion models to enhance natural image classification (He et al., 2023; Sarıyıldız et al., 2023; Trabucco et al., 2023; Azizi et al., 2023), segmentation (Sankaranarayanan et al., 2018; Wu et al., 2023b; Yang et al., 2024), and detection (Feng et al., 2024). Recently, Singh et al. (2024) performed an in-depth analysis of models trained with synthetic data across various robustness measures, and verified that they achieved good suitability in real-world settings.
To address data scarcity and promote diagnostic performance, the intelligent medical field has raised increasing interest in exploring diffusion-based generative augmentation for aiding different downstream diagnostic tasks (Shang et al., 2023; Farooq et al., 2024; Ktena et al., 2024; Zhang et al., 2024c; Dorjsembe et al., 2024b; Li et al., 2024b). Luo et al. (2024) marked the first focus on comprehensively analyzing the impact of their designed uncertainty-guided diffusion models on downstream diagnostic tasks. Nevertheless, current studies have two main limitations:
@@ 3.1 Model Overview
Fig. 3 shows the pipeline of using our proposed controllable generative augmentation framework to aid medical sequence classification. Our augmentation system is composed of two key components, including a sequence generator and a noisy synthetic data filter. First, the sequence generator perceives multiple semantic (including class label, descriptive text, and image prior) and sequential conditions (i.e., motion field) to produce customized and diverse synthetic databases. Then, the noisy synthetic data filter is employed to effectively suppress unsatisfactory generated sequences, resulting in reliability-enhanced and coherent synthetic databases for downstream diagnosis. In the application stage, the quality-controlled synthetic sequences and real ones work together to improve the performance of arbitrary classifiers. To formulate customized, high-fidelity, and coherent synthetic databases to boost classification, we design the whole framework from model-centric, i.e., sequence generator (Secs. 3.2, 3.3, and 3.4) and data-centric, i.e., data filter (Sec. 3.5) perspectives.
@@ 3.2 Basic Architecture of Sequence Generator
Fig. 4 shows the pipeline of our proposed sequence generator. It supports customized and high-quality medical sequence generation via multimodal conditions, using a two-stage training scheme. In the pretraining stage, it attends to high-fidelity visual features learning for controllable image synthesis. While in the finetuning stage, the domain-specific visual knowledge acquired from the previous stage is reused and focuses on sequential patterns modeling for customized sequence synthesis. During inference, given a single or combination of multimodal conditions as control signal inputs, high-quality and steerable sequence generation from Gaussian noise can work. We then provide a short background of video diffusion models, followed by a detailed description of the basic architecture of our generator.
Preliminaries of Video DDPMs. Video DDPMs aim to learn complex distributions by iteratively denoising corrupted inputs using diffusion methods (Sohl-Dickstein et al., 2015; Ho et al., 2020). Given a video $\mathit{x}_{0}\sim\mathit{q}$ and timestep $\mathit{T}$, DDPMs first produce a sequence of noisy inputs via diffusion process $\mathit{q}(\mathit{x}_{t}|\mathit{x}_{0},\mathit{t}),\mathit{t}\in{1,2,...,\mathit{T}}$, which progressively adds Gaussian noise $\epsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$ to $\mathit{x}_{0}$. To ease the computational burden of traditional pixel-space training, video latent diffusion models (Blattmann et al., 2023) perform the diffusion process in latent space of a variational autoencoder (Esser et al., 2021) (VAE). The model is then trained to estimate the parameterized Gaussian transition $\mathit{p}(\mathit{x}_{t-1}|\mathit{x}_{t})$ via a denoising network $\theta$. Mathematically, the optimized objective can be a simplified variant of the variational lower bound:
[…]
@@
Factorized Learning of Visual Features and Sequential Patterns. Developing a generator with pure 3D architecture design is an intuitive scheme for sequence synthesis (Blattmann et al., 2023). However, due to the scarcity and complexity of medical data, learning visual features and sequential patterns at the same time is very challenging and imposes high training costs. To this end, we propose to factorize both learning by first pretraining a latent diffusion model (LDM) and then finetuning its sequence counterpart. In this way, the generator enables realistic sequence synthesis, while easing the training burden.
2D-to-3D Model Inflation. During pretraining, we develop a 2D UNet (Ronneberger et al., 2015) in LDM to predict noises for image synthesis. Referring to (Ho et al., 2022b), we extend the 2D UNet to a 3D version via an inflation scheme to build the sequence LDM. Specifically, all the spatial convolution layers are inflated to pseudo-3D counterparts by expanding the kernels at the sequential dimension (e.g., $3\times 3\rightarrow 1\times 3\times 3$ kernel). Besides, we perform sequential insertion by adding sequential attention (SA) layers (see Fig. 4). Fig. 5(a) shows the principle of the SA mechanism, where each patch queries to those at the same spatial position and across frames/slices. This design encourages the generator to model the sequential patterns, while not significantly altering visual feature distribution baked in the LDM (Zhang et al., 2024a). Overall, sequence LDM enables inheriting the rich visual concepts preserved in LDM and focusing on sequential pattern aggregation, making the model learning efficient.
Figure 4: Pipeline of our proposed sequence generator. The MedSAM image encoder (Huang et al., 2024b) is used for domain-specific image prior feature extraction. During inference, given a bank of customized conditions (e.g., descriptive text with underlined carotid plaque attributes, the class label mild, image prior, and motion field), our generator produces a real-domain style sequence that faithfully adheres to all specified conditions.